The concentrated sulfuric acid we use in the laboratory is 98.0 percent H2SO4 by mass. Calculate the molality and molarity of the acid solution. The density of the solution is 1.83g/mL


1
Expert's answer
2022-02-19T06:51:38-0500

1 L = 1000 mL of solution

Mass of solution = volume x density

"Mass of solution = 1000\times1.83 = 1830 g=1000×1.83=1830g"


Mass of "H_2SO_4 =\frac{ 98.0}{100} × 1830 = 1793.4 g"


"Moles of H2SO4 = \frac{mass}{molar mass of H2SO4}"


"\frac{ 1793.4}{98.08} = 18.3 moles"

"Molarity =\frac{ moles of H2SO4}{volume of solution in L}" Molarity = moles of H2SO4/volume of solution in L

"\frac{18.3}{1} = 18.3 M"

Mass of water "= 1830 - 1793.4 = 36.6 g = 0.0366 kg"

"Molality = \frac{moles of H2SO4}{mass of water in kg}"

"\frac{18.3}{0.0366}

= 500 mol/kg = 500M"